OpenStudy (anonymous):

directions: write an equation for the line that satisfies each of the given conditions in the slope-intercept form (y=mx+b) and standard form (ax+by=c) question: slope=1/2, passes through (6,4)

jimthompson5910 (jim_thompson5910):

y = mx+b 4 = (1/2)(6) + b 4 = 6/2 + b 4 = 3+b 4-3 = b 1 = b So m = 1/2 and b = 1 which means that the equation is \[\large y = \frac{1}{2}x+1\]

OpenStudy (anonymous):

thanks